OpenStudy (thesmartestderpalive):

Can anyone help? Solve the word problem. The formula d = rt gives the distance traveled in time t at rate r. A bicyclist rides at a constant rate of 15 miles per hour. How many miles will he travel in 3.5 hours? A. 4.3 miles B. 18.5 miles C. 23.3 miles D. 52.5 miles <-----?

OpenStudy (thesmartestderpalive):

I think D. Because 15 x 3.5 = 52.5MPH

OpenStudy (anonymous):

Yes, you have the formula:\[d=rt=15*3.5=52.5\]

OpenStudy (anonymous):

It is MPH x H, so it gives 52.5 miles, not MPH @TheSmartestDerpAlive

OpenStudy (thesmartestderpalive):

Oh, Okay thanks! Here is a medal for you're doing! :)

OpenStudy (thesmartestderpalive):

Just a little more? I just want you to check my working. to let me understand what i'm doing...?

OpenStudy (thesmartestderpalive):

Solve the word problem. A group of roommates decide to split the cost of a new sofa. If s is the cost of the sofa and n is the number of roommates, the formula c=s/n determines each person’s share of the cost. If the new sofa costs $955 and 4 roommates live in the house, what is each person’s share? A. $382.00 B. $238.75 C. $221.25 D. $95.50

OpenStudy (thesmartestderpalive):

I say B. Because the formula is C=s/n, so 955(Divided by)4=238.75

OpenStudy (anonymous):

That is correct :)

OpenStudy (thesmartestderpalive):

Thanks!

OpenStudy (thesmartestderpalive):

I already got 100% though! :D